OpenStudy (selcouth):

A trolley company charges a fixed amount plus a fee based on the distance traveled. The cost chart for the company is shown: Cost Chart Distance traveled (miles) (x) Cost (dollars) (y) 0 6 1 10 2 14

OpenStudy (selcouth):

y = x + 36 y = 3x + 36 y = x + 30 y = 3x + 30

OpenStudy (anonymous):

Add it and multiply it all.

OpenStudy (anonymous):

in order to do this problem you must add and then multiply.